# Pragmatic Inference for Moral Reasoning Acquisition: Generalization via Distributional Semantics

### v1

@@ 7 Conclusion
In this paper, we explore methods to enhance generalization in moral reasoning by addressing the gap between the distributional semantics captured by LLMs and the inherently pragmatic nature of morality. Grounded in moral foundations theory, we first propose two pragmatic inference methods: one for classifying moral foundations from a given Rule-of-Thumb, and another for making moral judgments based on a given moral foundation. Building on these, we introduce a unified pragmatic inference method that enables LLMs to automatically infer moral foundations from a given moral situation and generate corresponding moral judgments. Our experimental results and analyses demonstrate that the proposed pragmatic inference methods significantly improve moral reasoning performance, reduce perplexity compared to baseline approaches, and avoid reliance on shallow heuristics.
@@ Appendix A Appendix
A.1 Definition of Moral Foundations (MF)
There are six moral foundations. Care: wanting someone or something to be safe, healthy, happy. Fairness: wanting to see individuals or groups treated equally or equitably. Liberty: wanting people to be free to make their own decisions. Loyalty: wanting unity and seeing people keep promises or obligations to an in-group. Authority: wanting to respect social roles, duties, privacy, peace, and order. Sanctity: wanting people and things to be clean, pure, innocent, and holy.
A.2 Experimental Details
We apply supervised fine-tuning with a learning rate of 5e-5 and a batch size of 24. All experiments are conducted for 5 epochs using three random seeds (1, 2, and 3), and we report the optimal result across three random seeds.
 Template for fine-tuning LLMs is:
base: The rule-of-thumb judgment is {RoT}. The moral foundations underlying the rule-of-thumb are {moral foundations}.
base+: {Definition of moral foundations} The rule-of-thumb judgment is {RoT}. The moral foundations underlying the rule-of-thumb are {moral foundations}.
ours: {Definition of moral foundations} The rule-of-thumb judgment is {RoT}. ###Inference: {inference} the moral foundations underlying the rule-of-thumb are {moral foundations}.
A.3 Answer Example for MFC
Rule of Thumb: Don’t think that printing money can fix all of your problems.
Inference: (a) The rule of thumb includes a Judgment to an Action. The Judgment is “can fix all of your problems” and the associated Action is “printing money”. (b) The consequence of the Action is that it would lead to severe inflation, devaluing the currency and causing economic instability. (c) The Action is relevant to the moral foundations care, fairness, and sanctity because its consequence harms people’s economic safety and health, violates fairness by unjustly redistributing wealth, and corrupts the sanctity of a stable monetary system which should be pure and trustworthy.
A.4 Answer Example for Moral Judgment
Prompt: Why don’t we just print a bunch of money to pay off our massive world debt? Reply: Because even if you find a way to print the money, what happens when it’s printed? The value of that money then becomes less and less, so when you try to pay off your debt, that’s when things get ugly.
Inference: (1) The moral foundations care, fairness, sanctity refer to a framework for understanding human morality where care involves protecting others from harm, fairness involves justice and proportional reciprocity, and sanctity involves respecting purity and avoiding degradation. (2) The conclusion of the Reply is that printing money to pay off debt is a bad idea because it leads to a severe decrease in the currency’s value which causes significant economic problems. (3) The conclusion of the Reply upholds care by preventing the widespread harm of hyperinflation which would hurt the most vulnerable people in society. It upholds fairness by preventing an unjust solution that would erode the real value of savings and contracts. It upholds sanctity by respecting the integrity and stability of the economic system rather than degrading it through a reckless and impure shortcut.
A.5 Answer Example for MFI-Judgment
Prompt: Why Don’t video plays have a real time clock in the control bar? Reply: Because the control bar does not support a time clock. Instead it simply counts up time to the end.
Inference: (1) The conclusion of the Reply is that the control bar’s design is a deliberate choice to show elapsed time rather than a real-time clock. (2) The conclusion of the Reply are relevant to moral foundations loyalty because it explains the control bar’s function as a kept promise to its in-group of users, fulfilling its obligation to display time in a specific, consistent way. (3) The conclusion of the Reply obeys the moral foundation of loyalty. It upholds the control bar’s obligation to its users by providing a consistent and honest account of its function, demonstrating unity between its stated design and its actual behavior

### v2

@@ 7 Conclusion
In this paper, we explore methods to enhance generalization in moral reasoning by addressing the gap between the distributional semantics captured by LLMs and the inherently pragmatic nature of morality. Grounded in Moral Foundations Theory, we first propose two pragmatic inference methods: one for classifying moral foundations from a given Rule of Thumb, and another for making moral judgments based on a given moral foundation. Building on these, we introduce a unified pragmatic inference method that enables LLMs to automatically infer moral foundations from a given moral situation and generate corresponding moral judgments. Our experimental results and analyses demonstrate that the proposed pragmatic inference methods significantly improve moral reasoning performance, reduce perplexity compared to baseline approaches, and avoid reliance on shallow heuristics.
@@ Appendix A Appendix
A.1 Experimental Details
We apply supervised fine-tuning with a learning rate of 5e-5 and a batch size of 24. For mistral-7B, we leverage the LoRA setting with rank: 64, lora alpha: 16, lora dropout: 0.1. All experiments are conducted for 5 epochs using three random seeds (1, 2, and 3), and we report the optimal result across three random seeds.  Template for fine-tuning LLMs is: base: The rule-of-thumb judgment is {RoT}. The moral foundations underlying the rule-of-thumb are {moral foundations}. base+: {Definition of moral foundations} The rule-of-thumb judgment is {RoT}. The moral foundations underlying the rule-of-thumb are {moral foundations}. ours: {Definition of moral foundations} The rule-of-thumb judgment is {RoT}. ###Inference: {inference} the moral foundations underlying the rule-of-thumb are {moral foundations}.
Figure 5: Definition of Six Moral Foundations
Figure 6: Prompts for Direct LLMs Prompting
Figure 7: Prompts for Standard CoT Prompting